Obtain the corresponding potential energy function of the intermolecular force. 1. F(r) = -ar^-7 + br^-9. 2. at what value of 'r' will the potential energy function be minimum


1
Expert's answer
2021-06-07T09:34:59-0400

"F(r)=-ar^{-7}+br^{-9},"

"F'(r)=\frac{7a}{r^8}-\frac{9b}{r^{10}}=\frac{7ar^2-9b}{r^{10}},"

"F'(r)=0,"

"r_{min}=\sqrt{\frac{9b}{7a}}."